Sean has 1 less than twice the number of stickers that Jason has. Rick has 53 more stickers than Jason has. If Sean gives Jason 25 stickers. Sean and Jason will have the same number of stickers. What is the total number of stickers the 3 boys have?
Answer by Edwin McCravy(20086)


Sean has 1 less than twice the number of stickers that Jason has.
         S = 2J - 1


Rick has 53 more stickers than Jason has.

         R = J + 53


If Sean gives Jason 25 stickers, Sean and Jason will have the same number of
stickers.
        S - 25 = J + 25

 
Solve the system of those three equation and get 

     Sean has 101, Jason has 51, and Rick has 104.

Checking:

Sean has 101, which 1 less than twice the number of stickers that Jason has,
which is 51.  Indeed that's true because twice 51 is 102 and Sean has 1 less
or 101.


Rick has 104, which is 53 more stickers than the 51 Jason has. That's true
since 51 + 53 = 104.

If Sean, who has 101, gives Jason 25 stickers, then Sean will have 101-25 or 76
and Jason will have 51+25 or 76, and that will indeed be the same number of
stickers, 76. 

So the answers are correct.  All you have to do is solve the system.  If you
can't, then post again asking how.

Edwin